OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9

OpenStudy (anonymous):

u serious? u need help in this question?

OpenStudy (anonymous):

Take minus 4x over to the left side and take minus 2 over to the right side 5x-4x=-9-2 x=-11

OpenStudy (anonymous):

5x-4x=-9-2 x=-11